Determine the differential canation x2 da u+1

Help me to solve this
Answer #1

dy = x² da at =) y= 22 2+1 = ) st , 이라 +) azuri => put uz atl I 2 ar =) 2+1 (U-12 u du Expand : u2 + 1 - 2 g = 1- 2 + 1 (0